Please help me solve this problem.
The measure of an angle is 20 less than the measure of its supplement. Find the measure of the angle, the measure of its supplement and the measure of its complement.

Answer by stanbon(48516)

The measure of an angle is 20 less than the measure of its supplement. Find the measure of the angle, the measure of its supplement and the measure of its complement.
---------------
EQUATION:
x + x+20 = 180
2x+20=180
2x=160
x=80 (one of the angles)
x+20 = 100 (the other angle)
===============
Cheers,
Stan H.